OpenStudy (anonymous):

Find the minimum y-value of the vertex. f(x)=x^2+2x-1 and f(x)=x^2-5x Please show work I would like to know how you got the answer to better understand it by someone explaining it.

jimthompson5910 (jim_thompson5910):

the min y value occurs at the vertex, so do you know how to find the vertex?

OpenStudy (anonymous):

y = a(x – h)2 + k?

jimthompson5910 (jim_thompson5910):

that's vertex form

jimthompson5910 (jim_thompson5910):

do you know how to find the x coordinate of the vertex?

OpenStudy (anonymous):

apparently not, I'm given these forms and examples and have to figure out what i can from them :/

jimthompson5910 (jim_thompson5910):

the x coordinate of y = ax^2 + bx + c is x = -b/(2a)

jimthompson5910 (jim_thompson5910):

y = x^2+2x-1 is the same as y = 1x^2+2x-1 in this case, a = 1 and b = 2, so ... x = -b/(2a) x = -2/(2*1) x = -2/2 x = -1

jimthompson5910 (jim_thompson5910):

this means that the x coordinate of the vertex for x^2+2x-1 is x = -1

jimthompson5910 (jim_thompson5910):

you then plug that back into y = x^2+2x-1 to get the y coordinate of the vertex

ganeshie8 (ganeshie8):

another way is to convert it to vertex form \(y = a(x – h)^2 + k?\) vertex = (h, k)

ganeshie8 (ganeshie8):

f(x)=x^2+2x-1 = (x+1)^2 -1 - 1 = (x--1)^2 -2 compare this with vertex form vertex = (-1, -2)

OpenStudy (anonymous):

I got -25/4 for x^2+2x-1, correct?

jimthompson5910 (jim_thompson5910):

no unfortunately it isn't

OpenStudy (anonymous):

f(x)=x^2-5x x=-b/2a =(-)-5/2*1 =5/2 x=f(5/2) =(5/2)^2-5*(5/2) =-25/4?

jimthompson5910 (jim_thompson5910):

oh you're working on a different function altogether

jimthompson5910 (jim_thompson5910):

you got it correct